OpenStudy (anonymous):

I award metals!! What is the arc length when Θ = pi over 3 and the radius is 5 cm?

OpenStudy (anonymous):

10 pi over 3cm 16 pi over 3cm pi over 3cm 5 pi over 3cm

OpenStudy (ipwnbunnies):

\[arc length = r \Theta\] Where the angle, theta, is in radians.

OpenStudy (anonymous):

Can you help explain this though? I have no idea what radians are

OpenStudy (anonymous):

like how would I fill in that equation

OpenStudy (ipwnbunnies):

Radians are a unit of measurement for an angle. They usually include pi. The angle they give you is in radians, so you can use the formula.

OpenStudy (ipwnbunnies):

What. p_p Your radius is 5 cm. Your angle is π/3 radians. \[arc length = r \Theta = (5 cm)(\frac{\pi}{3}) = ?\]

OpenStudy (anonymous):

so it would be like 5*1.04 since pi over 3 is rounded to 1.04?

OpenStudy (ipwnbunnies):

You don't need to convert it to decimal form. None of the choices are in decimal form. Use the true value, multiply the ratios.

OpenStudy (anonymous):

okay... so would it be the last one?

OpenStudy (ipwnbunnies):

Mhm.

OpenStudy (anonymous):

yay! Okay! Thanks so much!!

OpenStudy (anonymous):

There's your medal :)

OpenStudy (anonymous):

Thanks for the medal @iPwnBunnies! Can I ask you for your help on another one?

OpenStudy (ipwnbunnies):

Sure.

OpenStudy (anonymous):

Thank you!! The function f(t) = 3 cos(pi over 6t) + 5 represents the tide in Blastic Sea. It has a maximum of 8 feet when time (t) is 0 and a minimum of 2 feet. The sea repeats this cycle every 12 hours. After nine hours, how high is the tide?

OpenStudy (anonymous):

12 feet 5 feet 4.5 feet 2.5 feet

OpenStudy (anonymous):

How do I do this one?